Leno Manufacturing Company prepared the following factory overhead cost budget for the Press Department for October of the current year, during which it expected to require 9,000 hours of productive capacity in the department: Variable overhead cost: Indirect factory labor $66,600 Power and light 4,230 Indirect materials 29,700 Total variable overhead cost $100,530 Fixed overhead cost: Supervisory salaries $35,190 Depreciation of plant and equipment 22,120 Insurance and property taxes 14,070 Total fixed overhead cost 71,380 Total factory overhead cost $171,910 Assuming that the estimated costs for November are the same as for October, prepare a flexible factory overhead cost budget for the Press Department for November for 7,000, 9,000, and 11,000 hours of production. Round your interim computations to the nearest cent, if required. Enter all amounts as positive numbers.
Answer #1

Flexible factory overhead cost budget :-

7,000 hours 9,000 hours 11,000 hours
Variant overhead cost:
Indirect factory labor cost 7,000 x 7.40 = 51,800 66,600 11,000 x 7.40 = 81,400
Power and light 7,000 x 0.47 = 3,290 4,230 11,000 x 0.47 = 5,170
Indirect material 7,000 x 3.30 = 23,100 29,700 11,000 x 3.30 = 36,300
Total variable cost 78,190 100,530 122,870
Fixed overhead cost:
Supervisory salary 35,190 35,190 35,190
Depreciation 22,120 22,120 22,120
Insurance and property tax 14,070 14,070 14,070
Total fixed overhead 71,380 71,380 71,380
Total Factory overhead $149,570 $171,910 $194,250

Indirect factory labor cost per hour = 66,600/9,000 = 7.40

Power and light cost per hour = 4,230/9,000 = 0.47

Indirect Material cost per hour = 29,700/9,000 = 3.30